OpenStudy (patrickweewoos):

Natalie has to list numbers that are greater than 1. Which of these expressions represent numbers greater than 1?

OpenStudy (studygurl14):

x > 1

OpenStudy (anonymous):

z >1. Z must be greater.

OpenStudy (patrickweewoos):

A.)8-^2 Divided by 8-^3 B.)4^4 x 4-^3 C.)3^4 divided by 3^66 D.)7^3 x 7-^5 E.)2^8 Divided by 2^5

OpenStudy (valpey):

Is this the correct translation?: A: \(\frac{8^{-2}}{8^{-3}} = 8^{?}\) B: \(4^{4}*4^{-3} = 8^{?}\) C: \(\frac{8^{-2}}{8^{-3}} = 8^{?}\) D: \(7^{3}*7^{-5} = 7^{?}\) E: \(\frac{2^{8}}{2^{5}} = 2^{?}\)

OpenStudy (valpey):

Of course for B: I mean: \(=4^{?}\)